OpenStudy (anonymous):

HELP! 2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 (3x+6 8 ) = 24 1 (7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (phi):

is the original equation \[ \frac{3x+6}{8}= \frac{7x-1}{6} \]?

OpenStudy (phi):

if so, replace "x" with 1, and see if the equation is true (left side equals right side) if it is, x=1 is the correct answer.

OpenStudy (anonymous):

Yes that was the original equation. And it doesn't add up to the same

OpenStudy (phi):

to find the correct answer, we can multiply both sides by 2 to get \[\cancel{2} \frac{3x+6}{\cancel{8}4}= \cancel{2}\frac{7x-1}{\cancel{6}3} \] now "cross multiply to get to this step \[ 3(3x+6) = 4(7x−1) \] Max got this far, and then distributed the 3 on the left side, and the 4 on the right side: \[ 9x+18 = 28x -4 \] (max made a mistake and got 9x+18 = 28x−1 ) can you finish?

OpenStudy (anonymous):

Yes I think I got it now! Thank you :)